OpenStudy (anonymous):

Antiderivative (x-3)e^x

OpenStudy (cggurumanjunath):

use integration by parts.

OpenStudy (cggurumanjunath):

u= (x-3) v= e^x

OpenStudy (cggurumanjunath):